OpenStudy (anonymous):

A solid weighs 700 N. Calculate the pressure exerted by the solid on a square surface of side 20cm

OpenStudy (schrodingers_cat):

P = F/A

OpenStudy (anonymous):

P=F/A=700/0.04=17500 pa